OpenStudy (anonymous):

solve for y: (x+2)^2+(y-3)^2=25

jimthompson5910 (jim_thompson5910):

\[\large (x+2)^2+(y-3)^2=25\] \[\large (y-3)^2=25-(x+2)^2\] \[\large y-3=\pm\sqrt{25-(x+2)^2}\] \[\large y=\pm\sqrt{25-(x+2)^2}+3\]

OpenStudy (anonymous):

really? this is a circle with center (-2,3) and radius 5

OpenStudy (anonymous):

will get two solutions

OpenStudy (anonymous):

oh jimthompson has it worked out

jimthompson5910 (jim_thompson5910):

So you basically have two equations \[\large y=3+\sqrt{25-(x+2)^2}\] \[\large y=3-\sqrt{25-(x+2)^2}\]